# Efficient classical algorithms for linear optical circuits

### v2

@@ Abstract
Corollary 3 (Phase shifter expectation value).
Phase shifter expectation value of a linear optical circuit can be efficiently approximated for an arbitrary product input state.
Therefore, our algorithm enables efficient approximation of the grouped probabilities in Eq. (9) for arbitrary input states and thus solves a generalized molecular vibronic spectra problem for arbitrary input states, an open problem posed in Ref. [36].
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to near-Clifford circuits $\hat{U}_{\text{NC}}$ [55, 56], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary input states. Note that exact simulation of such a circuit is hard, similar to the case of boson sampling [75]. Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
@@
Consider an $M$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(M))$ and an operators $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(M)\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$.
@@
$\langle\psi|\hat{U}^{\dagger}_{\text{C}}\hat{O}\hat{U}_{\text{C}}|\psi\rangle=\frac{1}{2^{l}}\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\chi_{\hat{O}_{A}}(\bm{a})\chi_{\hat{\rho}_{A}}(\bm{a})$ (10)
$=\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\frac{|\chi_{\hat{O}_{A}}(\bm{a})|^{2}}{2^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})}{\chi^{*}_{\hat{O}_{A}}(\bm{a})}.$ (11)
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}$, the median-of-means estimator gives the desired error presented in the theorem.
Furthermore, our method is still valid for the additional logarithmic number of $T$-gates because a single Pauli operator is transformed into the sum of at most two Pauli operators after passing through each $T$-gate, that is, $\hat{T}^{\dagger}\hat{P}_{0}\hat{T}=\hat{P}_{0},~\hat{T}^{\dagger}\hat{P}_{1}\hat{T}=(\hat{P}_{1}+\hat{P}_{2})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{2}\hat{T}=(\hat{P}_{2}-\hat{P}_{1})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{3}\hat{T}=\hat{P}_{3}$. If there are $t$ of $T$-gates in a near-Clifford circuit $\hat{U}_{\text{NC}}$, we have at most $M2^{t}$ Pauli operators to track the transformations. Therefore, if $t=O(\log M)$, we can handle all the relevant Pauli operators in $O(\text{poly}(M))$ time. A detailed proof is given in SM Sec. S6 [57]. ∎
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(M))$ and marginal probabilities of a near-Clifford circuit with an arbitrary input state can be efficiently approximated. Thus, for approximating the expectation values of near-Clifford circuits, the magic of input states does not increase the complexity [76]. Hence, our result is a generalization of previous results: the expectation values of a near-Clifford circuit for $|\psi\rangle=|0^{M}\rangle$ or for Pauli observables [55, 56] and (marginal) output probabilities of a Clifford circuit [77, 66].
Discussion.— We presented an efficient classical algorithm for estimating expectation values of observables in linear optical circuits. Using this, we showed that one can efficiently approximate (marginal) output probabilities of boson sampling with arbitrary input states, which implies that the simulation of sparse boson sampling with arbitrary input states is easy. We also generalized Gurvits’s algorithm to more general amplitudes that appear in linear optical circuits, which allows us to efficiently approximate a hafnian. Furthermore, using this classical algorithm, we solved an open problem in Ref. [36], namely, a generalized molecular vibronic spectra problem for arbitrary input states.
@@
Also, we have focused on computing grouped probabilities of boson sampling with a certain pattern in the molecular vibronic spectra problem, which is defined by the linear relation $\Omega-\bm{\omega}\cdot\bm{m}$ in Eq. (9). However, our method does not straightforwardly generalize to a more complicated pattern of groups, which still opens a possibility of quantum advantage. Thus, the complexity of approximating general grouped boson sampling output probabilities is another open question.

### v3

@@ Abstract
Corollary 3 (Phase-shifter expectation value).
Phase-shifter expectation value of a linear-optical circuit can be efficiently approximated for an arbitrary product input state.
Therefore, our algorithm enables efficient approximation of the grouped probabilities in Eq. (7) for arbitrary product input states and thus solves a generalized MVS problem for arbitrary product input states, an open problem posed in Ref. [39].
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to qubit system, especially near-Clifford circuits $\hat{U}_{\text{NC}}$ [65, 66], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary product input states. It is known that both exact and approximate classical simulations of such circuits are computationally hard, assuming the non-collapse of the polynomial hierarchy [88]. Consequently, to investigate routes toward practical quantum advantage, our method provides a direct additive-error estimation scheme for the expectation value of an observable, $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ that does not rely on sampling from the output distribution, thereby bypassing the main bottleneck that underlies the known sampling hardness results.
Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear-optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
@@
Consider an $n$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(n))$ and an operators $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive-error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(n)\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]\log(1/\delta)/\epsilon^{2})$.
@@
$\langle\psi|\hat{U}^{\dagger}_{\text{C}}\hat{O}\hat{U}_{\text{C}}|\psi=\sum_{\bm{a}\in\{0,1,2,3\}^{l}}\frac{|\chi_{\hat{O}_{A}}(\bm{a})|^{2}}{2^{l}\|\hat{O}_{A}\|_{2}^{2}}\frac{\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})}{\chi^{*}_{\hat{O}_{A}}(\bm{a})}.$ (8)
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\mbox{Tr}[\hat{\rho}_{A}^{2}]$, the median-of-means estimator gives the desired error presented in the theorem.
Furthermore, our method is still valid for the additional logarithmic number of $T$-gates because a single Pauli operator is transformed into the sum of at most two Pauli operators after passing through each $T$-gate, that is, $\hat{T}^{\dagger}\hat{P}_{0}\hat{T}=\hat{P}_{0},~\hat{T}^{\dagger}\hat{P}_{1}\hat{T}=(\hat{P}_{1}+\hat{P}_{2})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{2}\hat{T}=(\hat{P}_{2}-\hat{P}_{1})/\sqrt{2},~\hat{T}^{\dagger}\hat{P}_{3}\hat{T}=\hat{P}_{3}$. If there are $t$ of $T$-gates in a near-Clifford circuit $\hat{U}_{\text{NC}}$, we have at most $n2^{t}$ Pauli operators to track the transformations. Therefore, if $t=O(\log n)$, we can handle all the relevant Pauli operators in $O(\text{poly}(n))$ time. A detailed proof is given in SM Sec. S7 [40]. ∎
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(n))$ of a near-Clifford circuit with an arbitrary product input state can be efficiently approximated. Especially for rank-1 projective measurements, this result implies that the magic resources of both the input state and measurement do not increase the complexity of additive-error-approximation. Note that our result far extends the previous results: the expectation values of a near-Clifford circuit with stabilizer input or for Pauli observables [65, 66], and output probabilities in the computational basis [89, 66, 82]. (See SM Sec. S10 for more details [40].)
Discussion.— We presented a classical algorithm for estimating expectation values of observables in linear-optical circuits. Using this, we showed that one can efficiently approximate (marginal) output probabilities of boson sampling with arbitrary product input states, which implies that the simulation of sparse boson sampling with arbitrary product input states is easy. We also provided an efficient classical algorithm for the general transition amplitude in a linear-optical circuit, which allows us to approximate the hafnian of any complex symmetric matrix. Furthermore, using this classical algorithm, we solved an open problem in Ref. [39], namely, a generalized MVS problem for arbitrary product input states.
@@
Also, we have focused on computing grouped probabilities of boson sampling with a certain pattern in the MVS problem, which is defined by the linear relation $\Omega-\bm{\omega}\cdot\bm{m}$ in Eq. (7). However, our method does not straightforwardly generalize to a more complicated pattern of groups, which still opens a possibility of quantum advantage. Thus, the complexity of approximating general grouped boson sampling output probabilities is another open question.